# Automatic Depression Assessment using Machine Learning: A Comprehensive Survey

### v1

@@ 4.6. Language/text-based depression assessment
Language/textual data (e.g., conversations, social media posts, and clinical notes) provide semantic meaningful verbal information that can directly and individually identify human depression status (Coppersmith et al., 2015).
4.6.1. Challenges.
@@
As discussed in Sec. 2.2.5, the key challenge for language/text-based ADA is the cross-language analysis, as some depression biomarkers are language-dependent. Consequently, Chen et al. (Chen et al., 2023) developed separate models (i.e., Chinese BERT and English BERT models) to generate sentence-level embeddings from texts of different domains. Another common challenge is to account for multi-scale (word-level, sentence-level and topic level) contexts when analysing language-based depression cues, where a frequently employed solution is to extract and combine multi-scale languages features (De Choudhury et al., 2013b, a; Tadesse et al., 2019; Shen et al., 2017), or extract word/sentence-level features by specifically considering conversation/text topics (Lam et al., 2019; Trotzek et al., 2018; Gong and Poellabauer, 2017). We detailed review these text-based ADA approaches in the subsection below.
4.6.2. Hand-crafted approaches:
@@
Hand-crafted language features for ADA can be categorised into three primary scales: lexical (word-level), syntactic (fragmented sentence-level), and semantic (topic-level). Each scale provides unique insights into language use and psychological depression indicators. Specifically, lexical features are extracted from individual words to describe their linguistic or emotional attributes (e.g., pronouns, verbs, negative emotion words, and anger words), where Pennebaker’s Linguistic Inquiry and Word Count (LIWC) (Pennebaker et al., 2001) is a prevalent feature extraction strategy for text-based ADA (Resnik et al., 2013; De Choudhury et al., 2013b; Shen et al., 2017; Coppersmith et al., 2015; Nguyen et al., 2014; Tadesse et al., 2019; De Choudhury et al., 2014; Gratch et al., 2014). Additionally, more tailored lexicons such as word count, pronoun usage, and other stylistic markers (detailed discussed in Sec. 2.2.5) that are critical for depression assessment also have been widely utilised (De Choudhury et al., 2013b). Meanwhile, syntactic features focus on the structural patterns of language at the fragmented sentence-level, as fragmented sentences expressed by depressed individuals often reflect reduced coherence and cognitive difficulty (Resnik et al., 2015). These include the number of replies, shares, or interactions indicating individuals’ engagement in social media contexts (De Choudhury et al., 2013b), i.e., reduced social interaction or lower mention rate have been proven to be strong depression indicators (De Choudhury et al., 2013b; Tsugawa et al., 2015). Besides, semantic features operate at a more broader level to capture semantics and themes provided by languages. A widely-used depression semantic feature is extracted through Latent Dirichlet Allocation (LDA) (Blei et al., 2003), which identifies latent topics reflecting patterns of language use and thematic areas associated with depression (e.g., expressions of loneliness, hopelessness and mental health concerns (Resnik et al., 2013; Shen et al., 2017; Tsugawa et al., 2015; Nguyen et al., 2014; Tadesse et al., 2019; Resnik et al., 2015)).
@@
As these three types of linguistic features capture complementary depression-related information at different scales, they have been frequently combined for ADA, with a large part of them being applied to social media data. For example, Choudhury et al. (De Choudhury et al., 2013b) analyse Twitter data from the year preceding depression onset by extracting lexical features (e.g., LIWC, as well as terms related to anxiety, insomnia and antidepressant mentions), syntactic features (e.g., reply proportion and question-focused posts), and semantic features (e.g., post frequency, shares, posting time and social graph metrics) characterizing social interactions. Finally, an SVM integrates these features for ADA. Its extension (De Choudhury et al., 2013a) further monitors population-level depression trends in real-time, additionally analysing unigrams (an n-gram consisting of a single item from a sequence, such as a single word) and bigrams (a pair of consecutive written units such as letters, syllables, or words) from Twitter posts to account for the general language use, and then applies a PCA and SVM for more accurate ADA. Shen et al. (Shen et al., 2017) construct a publicly available Twitter dataset, from which three types of features including LIWC emotional linguistic features, semantic LDA topic features, and social interaction features such as the numbers of tweets, followers and followings, are extracted, which are then jointly encoded into a sparse representation and processed by a regularised LR model for ADA. Based on LIWC features and user engagement metrics, Coppersmith et al. (Coppersmith et al., 2015) incorporate open-vocabulary methods to specifically character n-gram language models, for capturing depression condition-specific character-level linguistic features. Besides Twitter, texts appeared in other social media channels have also been utilised for ADA. Tadesse et al. (Tadesse et al., 2019) develop their approach on a Reddit posts dataset annotated with binary depression labels. It adopts LIWC, LDA, and n-gram modeling to extract features related to correlation significance, hidden topics and word frequencies. The three textual feature sets are then concatenated and fed to various ML classifiers for ADA. Apart from social media contents, Resnik et al. (Resnik et al., 2013) apply LIWC and LDA to extract lexical and semantic features from student essays. It categorises words into psychological groups and employs topic modeling to enhance depression feature extraction. Their subsequent work (Resnik et al., 2015) extends LDA as a supervised nested LDA (SNLDA) to specifically extract depression-related semantic contents for ADA. Based on LiveJournal, Nguyen et al. (Nguyen et al., 2014) extracts LIWC features, LDA features, affective features, and mood tags selected by users for their posts, where a regularised regression model is employed to refine features and make binary depression decisions. Sun et al. (Sun et al., 2017) analyse DAIC-WOZ interview transcripts to define six depression-related attributes (e.g., sleep quality, PTSD diagnosis, treatment history, personal preferences, emotions). They then identify five most depression-related attributes, and fuse their predictions using a cascade strategy with a RF model to produce the final depression severity. Tsugawa et al. (Tsugawa et al., 2015) employ bag-of-words to extract word-level linguistic features (e.g., emotional tone, word frequency), while LDA derives semantic content features. Additionally, social interaction metrics (retweet rate, mention rate and URL usage) are further extracted. These features are concatenated into a single vector and processed via SVM for ADA.
4.6.3. Deep learning approaches:
@@
A typical DL-based language ADA solution (Shen et al., 2022) adopts ELMo to extract sentence-level embeddings and then feed them to a DL network consisting of two BiLSTM layers and an attention layer. This results in a contextualised text representation that emphasises the most depression-relevant sentences, which is then concatenated with audio features via a FC layer for depression assessment. Hong et al. (Hong et al., 2022) propose a schema-based GNN which schematises underlying associations among words (i.e., node embeddings), aiming to capture depression symptoms within the entire transcript. Alternatively, DL-based approaches also frequently model semantics related to question topics and responses of the given language/textual data. Lam et al. (Lam et al., 2019) identify every topic from transcripts and generate augmented data by shuffling topic segments within selected combinations. The augmented texts are then used to train a transformer for deep depression-related textual feature extraction. During inference, the extracted features are integrated with features of other modalities via a dense classification layer for ADA. Based on a Chinese dataset recorded individuals’ conversation with physician, Ye et al. (Ye et al., 2021) employ Word2Vec for initial word embedding extraction, and a one-hot Transformer for deep feature extraction. These features are finally concatenated with features from audio modality for ADA. Based on posts and comments from Reddit, Trotzek et al. (Trotzek et al., 2018) propose two approaches. The first applies LIWC to extract hand-crafted features capturing linguistic and psychological characteristics, which are then fed into a LR classifier. The second applies fastText and GloVe to learn word-level embeddings to model semantic relationships in the text, followed by a CNN to learn depression-specific cues. Alhanai et al. (Al Hanai et al., 2018) apply Doc2Vec (GENSIM, 2024) to generate an embedding of every textual responses, which further understands the varying prognostic relevance to depression from them. These features are finally processed by LSTMs and FC layers for ADA.
4.7. Multi-modal automatic depression assessment
@@
Since different human behaviour modalities could provide complementary cues for depression assessment (Liu et al., 2024; Cheong et al., 2024c; Zheng et al., 2023), a large number of approaches attempted to leverage multi-modal human behaviours for ADA.
4.7.1. Challenges.
@@
The most common challenge unique to multi-modal ADA systems is when to fuse multi-modal features. Typical solutions include combining frame/segment-level multi-modal features (Uddin et al., 2022; Yoon et al., 2022; Zhang, [n. d.]), fusion of multi-modal features at multiple temporal scales (Hu et al., 2024; Shen et al., 2024), fusion of clip-level predictions achieved for all modalities (Senoussaoui et al., 2014; Valstar et al., 2016; Ringeval et al., 2017; Williamson et al., 2016; Pampouchidou et al., 2016), and fusion of multi-modal clip-level representations (Dibeklioğlu et al., 2018; Scherer et al., 2013a; Alghowinem et al., 2016; Tao et al., 2024; Niu et al., 2020; Pampouchidou et al., 2016; Gupta et al., 2014; Nasir et al., 2016; Zhang et al., 2024a; Wei et al., 2022; Zhang et al., 2024c). How to adaptively combine multi-modal features/predictions is also a key challenge, where average of multi-modal predictions (Valstar et al., 2013, 2014; Senoussaoui et al., 2014; Valstar et al., 2016; Ringeval et al., 2017), multi-modal feature concatenation (Dibeklioğlu et al., 2018; Scherer et al., 2013a; Pampouchidou et al., 2016; Gupta et al., 2014; Nasir et al., 2016; Yang et al., 2018; Rodrigues Makiuchi et al., 2019), Gaussian Staircase Model (Williamson et al., 2016), cross-modal attentions (Niu et al., 2020; Zhang et al., 2024a, c), graph (Shen et al., 2024) and other DL operations (Uddin et al., 2022; Cheong et al., 2024b; Hu et al., 2024; Wei et al., 2022; Lin et al., 2021) have been frequently utilised.
4.7.2. Audio-visual approaches
@@
Hand-crafted approaches: Early multi-modal ADA solutions also frequently rely on different combinations of hand-crafted features and traditional ML predictors. The baselines of AVEC 2013 and AVEC 2014 challenges (Valstar et al., 2013, 2014) extract various hand-crafted image-level static and spatio-temporal facial descriptors (discussed in Sec. 4.3), as well as apply openSMILE (Eyben et al., 2010) to extract frame-level LLD audio features. Then, clip-level audio and visual depression predictions are obtained by separately feeding clip-level audio and visual representations (i.e., mean vectors across all frame-level features) to SVR models, followed by a decision-level fusion. Senoussaoui et al. (Senoussaoui et al., 2014) extract i-vector representations from frame-level MFCC features as the clip-level audio features while averaging all LGBP-TOP features summarising short-term facial behaviours as the clip-level visual representation, based on which unimodal audio/facial depression occurrence or severity are individually predicted via traditional ML predictors (e.g., SVM), where the decision-level fusion is conducted to make the final clip-level multi-modal prediction. Dibekliouglu et al. (Dibeklioğlu et al., 2018) first apply Stacked Denoising Autoencoders (SDAE) to encode all frame-level facial expression and head motions that are related to depression status, and then summarise them as a fixed-size clip-level representation via IFV coding and Compact Dynamic Feature Set. Meanwhile, clip-level audio features including fundamental frequency, switching pause duration, Maxima Dispersion Quotient and Peak Slope are also extracted. These clip-level features are finally enhanced by a feature selection, and combined to be processed by LR for ADA. Scherer et al. (Scherer et al., 2013a) extract three clip-level audio features, including differentiated glottal flow, quasi-open quotient and dyadic wavelet transform feature. These features are concatenated with two clip-level facial features: emotional variability computed from all frame-level discrete facial expressions and motor variability describing hand gesturing and head movements, which are finally fed to SVM for depression classification. Instead of directly using original face videos, the baselines of AVEC 2016 and AVEC 2017 challenges (Valstar et al., 2016; Ringeval et al., 2017) develop their ADA approaches based on audio behaviours and anonymous frame-level facial primitives to avoid ethical/privacy issues (discussed in Sec. 4.3), where hand-crafted prosodic, voice quality and spectral features are extracted to summarise short-term audio behaviours. Then, depression severity is predicted frame-wisely using RF regressor, and the unimodal clip-level audio and visual depression predictions are individually achieved by averaging all frame-level predictions, which are further averaged to achieve clip-level multi-modal ADA. Alghowinem et al. (Alghowinem et al., 2016) compute clip-level statistical features from not only audio behaviours but also frame-level eye activity and head poses, which are concatenated and undergo various feature selection processes, with the selected feature being fed to SVM for depression classification.
@@
Deep learning approaches: A common pipeline for DL-based audio-visual ADA starts with learning short-term/frame-level behaviours, which are then encoded as clip-level representations to make the clip-level multi-modal depression assessment. For example, Uddin et al. (Uddin et al., 2022) first split the audio-visual data into a set of fixed-length segments and then feed them to a spatio-temporal network for depression-related segment-level multi-modal feature learning, identifying behaviours that are mostly contributed to depression assessment. These segment-level features are finally aggregated as clip-level representation for ADA. Tao et al. (Tao et al., 2024) separately encode the audio and visual segment-level features as clip-level representations via a temporal attentive pooling, which are then fused via a multi-modal factorised bilinear pooling to provide the final prediction. Niu et al. (Niu et al., 2020) propose a spatio-temporal attention network to separately learn short-term facial and audio depression cues, and then apply the eigen-evolution pooling to aggregate all short-term audio-visual cues as the clip-level representation. Finally, an attention module and a SVR complementarily combines clip-level multi-modal representations for ADA. Similarly, Yoon et al. (Yoon et al., 2022) employ a two-stream transformer to learn clip-level facial and audio depression representations from all frame-level facial landmarks and audio LLDs, respectively. Then, a cross-attention module is adopted to capture the relationship between audio and visual features to learn a clip-level multi-modal representation for ADA. Focusing on fair ADA, Cheong et al. (Cheong et al., 2024b) propose a causal multi-modal framework for debiasing multi-modal behaviours expressed by individuals with different backgrounds where audio and visual features are adaptively fused via a novel referee-based individual fairness guided fusion mechanism to make the multi-modal depression prediction.
@@
Hybrid approaches: Some approaches also manage to leverage both domain knowledge provided by hand-crafted features and task-specific feature learning capability of the DL models for ADA. The AVEC 2019 challenge baseline (Ringeval et al., 2019) summarise various audio LLDs and facial behaviour primitives (e.g., AUs) of the given audio-visual clip. Jan et al. (Jan et al., 2017) extract multiple hand-crafted frame-level facial features (LBP, EOH and LPQ), with a CNN learning depression-specific frame-level facial features. Meanwhile, spectral audio LLDs and MFCC features are extracted from every audio segment. The dynamics of these audio and visual features are then modelled by the feature dynamic history histogram (FDHH) for clip-level ADA. Hu et al. (Hu et al., 2024) propose a two-stream CNN-LSTM network which separately capture short-term dynamics/features from all frame-level facial primitives (i.e., facial landmarks, gaze and AUs), and audio log-mel spectrograms. Then, a spatio-temporal attention pooling is proposed to obtain long-term audio-visual representations, which are fused at multiple temporal scales via a bridge fusion to output the clip-level multi-modal depression prediction.
4.7.3. Audio-visual-language approaches

### v2

@@ 4.6 Language/text-based depression assessment
Language/textual data (e.g., conversations, social media posts, and clinical notes) provide semantic meaningful verbal information that can directly and individually identify human depression status [289].
4.6.1 Challenges and solutions
@@
As discussed in Sec. 2.2.5, the key challenge for language/text-based ADA is that some verbal depression biomarkers are language-dependent. Consequently, Chen et al. [271] developed separate models (i.e., Chinese BERT and English BERT models) to generate sentence-level embeddings from texts of different domains. Another common challenge is utilising multi-scale (word-level, sentence-level and topic level) contexts when analysing depression-related language cues, where a frequently employed solution is to extract and combine multi-scale languages features [133, 135, 234, 233], or extract word/sentence-level features by specifically considering conversation/text topics [272, 235, 36]. We detailed review language/text-based ADA approaches in the subsection below.
4.6.2 Hand-crafted approaches
@@
Hand-crafted language features for ADA can be categorised into three primary scales: lexical (word-level), syntactic (fragmented sentence-level), and semantic (topic-level). Each scale provides unique insights into language use and psychological depression indicators. Specifically, lexical features are extracted from individual words to describe their linguistic or emotional attributes (e.g., pronouns, verbs, negative emotion words, and anger words), where Pennebaker’s Linguistic Inquiry and Word Count (LIWC) [22] is a prevalent feature extraction strategy for text-based ADA [237, 133, 233, 289, 232, 234, 290, 64]. Additionally, more tailored lexicons such as word count, pronoun usage, and other stylistic markers (detailed discussed in Sec. 2.2.5) that are critical for depression assessment also have been widely utilised [133]. Meanwhile, syntactic features focus on the structural patterns of language at the fragmented sentence-level, as fragmented sentences expressed by depressed individuals often reflect reduced coherence and cognitive difficulty [137]. These include the number of replies, shares, or interactions indicating individuals’ engagement in social media contexts [133], i.e., reduced social interaction or lower mention rate have been proven to be strong depression indicators [133, 128]. Besides, semantic features operate at a more broader level to capture semantics and themes provided by languages. A widely-used depression semantic feature is extracted through Latent Dirichlet Allocation (LDA) [291], which identifies latent topics reflecting patterns of language use and thematic areas associated with depression (e.g., expressions of loneliness, hopelessness and mental health concerns [237, 233, 128, 232, 234, 137]).
@@
As these three types of linguistic features capture complementary depression-related information at different scales, they have been frequently combined for ADA, with a large part of them being applied to social media data. For example, Choudhury et al. [133] analyse Twitter data from the year preceding depression onset by extracting lexical features (e.g., LIWC, as well as terms related to anxiety, insomnia and antidepressant mentions), syntactic features (e.g., reply proportion and question-focused posts), and semantic features (e.g., post frequency, shares, posting time and social graph metrics) characterizing social interactions. Finally, an SVM integrates these features for ADA. Its extension [135] further monitors population-level depression trends in real-time, additionally analysing unigrams (an n-gram consisting of a single item from a sequence, such as a single word) and bigrams (a pair of consecutive written units such as letters, syllables, or words) from Twitter posts to account for the general language use, and then applies a PCA and SVM for more accurate ADA. Shen et al. [233] construct a publicly available Twitter dataset, from which three types of features including LIWC emotional linguistic features, semantic LDA topic features, and social interaction features such as the numbers of tweets, followers and followings, are extracted, which are then jointly encoded into a sparse representation and processed by a regularised LR model for ADA. Based on LIWC features and user engagement metrics, Coppersmith et al. [289] incorporate open-vocabulary methods to specifically character n-gram language models, for capturing depression condition-specific character-level linguistic features. Besides Twitter, texts appeared in other social media channels have also been utilised for ADA. Tadesse et al. [234] develop their approach on a Reddit posts dataset annotated with binary depression labels. It adopts LIWC, LDA, and n-gram modeling to extract features related to correlation significance, hidden topics and word frequencies. The three textual feature sets are then concatenated and fed to various ML classifiers for ADA. Apart from social media contents, Resnik et al. [237] apply LIWC and LDA to extract lexical and semantic features from student essays. It categorises words into psychological groups and employs topic modeling to enhance depression feature extraction. Their subsequent work [137] extends LDA as a supervised nested LDA (SNLDA) to specifically extract depression-related semantic contents for ADA. Based on LiveJournal, Nguyen et al. [232] extracts LIWC features, LDA features, affective features, and mood tags selected by users for their posts, where a regularised regression model is employed to refine features and make binary depression decisions. Sun et al. [222] analyse DAIC-WOZ interview transcripts to define six depression-related attributes (e.g., sleep quality, PTSD diagnosis, treatment history, personal preferences, emotions). They then identify five most depression-related attributes, and fuse their predictions using a cascade strategy with a RF model to produce the final depression severity. Tsugawa et al. [128] employ bag-of-words to extract word-level linguistic features (e.g., emotional tone, word frequency), while LDA derives semantic content features. Additionally, social interaction metrics (retweet rate, mention rate and URL usage) are further extracted. These features are concatenated into a single vector and processed via SVM for ADA.
4.6.3 Deep learning approaches
@@
A typical DL-based language ADA solution [69] adopts ELMo to extract sentence-level embeddings and then feed them to a DL network consisting of two BiLSTM layers and an attention layer. This results in a contextualised text representation that emphasises the most depression-relevant sentences, which is then concatenated with audio features via a FC layer for depression assessment. Hong et al. [257] propose a schema-based GNN which schematises underlying associations among words (i.e., node embeddings), aiming to capture depression symptoms within the entire transcript. Alternatively, DL-based approaches also frequently model semantics related to question topics and responses of the given language/textual data. Lam et al. [272] identify every topic from transcripts and generate augmented data by shuffling topic segments within selected combinations. The augmented texts are then used to train a transformer for deep depression-related textual feature extraction. During inference, the extracted features are integrated with features of other modalities via a dense classification layer for ADA. Based on a Chinese dataset recorded individuals’ conversation with physician, Ye et al. [243] employ Word2Vec for initial word embedding extraction, and a one-hot Transformer for deep feature extraction. These features are finally concatenated with features from audio modality for ADA. Based on posts and comments from Reddit, Trotzek et al. [235] propose two approaches. The first applies LIWC to extract hand-crafted features capturing linguistic and psychological characteristics, which are then fed into a LR classifier. The second applies fastText and GloVe to learn word-level embeddings to model semantic relationships in the text, followed by a CNN to learn depression-specific cues. Alhanai et al. [224] apply Doc2Vec [292] to generate an embedding of every textual responses, which further understands the varying prognostic relevance to depression from them. These features are finally processed by LSTMs and FC layers for ADA.
4.7 Multi-modal automatic depression assessment
@@
Since different human behaviour modalities could provide complementary cues for depression assessment [190, 249, 246], a large number of approaches attempted to leverage multi-modal human behaviours for ADA.
4.7.1 Challenges and solutions
@@
A common challenge unique to multi-modal ADA systems is when to fuse multi-modal features. Typical solutions include combining frame/segment-level multi-modal features [258, 68, 263], fusion of multi-modal features at multiple temporal scales [259, 266], fusion of clip-level predictions achieved for all modalities [240, 61, 62, 268, 25], and fusion of multi-modal clip-level representations [66, 269, 241, 248, 29, 25, 251, 239, 270, 245, 273]. How to adaptively combine multi-modal features/predictions is also a key challenge, where average of multi-modal predictions [59, 60, 240, 61, 62], multi-modal feature concatenation [66, 269, 25, 251, 239, 275, 37], Gaussian Staircase Model [268], cross-modal attentions [29, 270, 273], graph [266] and other DL operations [258, 244, 259, 245, 67] have been frequently utilised.
4.7.2 Audio-visual approaches
@@
Hand-crafted approaches: Early multi-modal ADA solutions also frequently rely on different combinations of hand-crafted features and traditional ML predictors. The baselines of AVEC 2013 and AVEC 2014 challenges [59, 60] extract various hand-crafted image-level static and spatio-temporal facial descriptors (discussed in Sec. 4.3), as well as apply openSMILE [293] to extract frame-level LLD audio features. Then, clip-level audio and visual depression predictions are obtained by separately feeding clip-level audio and visual representations (i.e., mean vectors across all frame-level features) to SVR models, followed by a decision-level fusion. Senoussaoui et al. [240] extract i-vector representations from frame-level MFCC features as the clip-level audio features while averaging all LGBP-TOP features summarising short-term facial behaviours as the clip-level visual representation, based on which unimodal audio/facial depression occurrence or severity are individually predicted via traditional ML predictors (e.g., SVM), where the decision-level fusion is conducted to make the final clip-level multi-modal prediction. Dibekliouglu et al. [66] first apply Stacked Denoising Autoencoders (SDAE) to encode all frame-level facial expression and head motions that are related to depression status, and then summarise them as a fixed-size clip-level representation via IFV coding and Compact Dynamic Feature Set. Meanwhile, clip-level audio features including fundamental frequency, switching pause duration, Maxima Dispersion Quotient and Peak Slope are also extracted. These clip-level features are finally enhanced by a feature selection, and combined to be processed by LR for ADA. Scherer et al. [269] extract three clip-level audio features, including differentiated glottal flow, quasi-open quotient and dyadic wavelet transform feature. These features are concatenated with two clip-level facial features: emotional variability computed from all frame-level discrete facial expressions and motor variability describing hand gesturing and head movements, which are finally fed to SVM for depression classification. Instead of directly using original face videos, the baselines of AVEC 2016 and AVEC 2017 challenges [61, 62] develop their ADA approaches based on audio behaviours and anonymous frame-level facial primitives to avoid ethical/privacy issues (discussed in Sec. 4.3), where hand-crafted prosodic, voice quality and spectral features are extracted to summarise short-term audio behaviours. Then, depression severity is predicted frame-wisely using RF regressor, and the unimodal clip-level audio and visual depression predictions are individually achieved by averaging all frame-level predictions, which are further averaged to achieve clip-level multi-modal ADA. Alghowinem et al. [241] compute clip-level statistical features from not only audio behaviours but also frame-level eye activity and head poses, which are concatenated and undergo various feature selection processes, with the selected feature being fed to SVM for depression classification.
@@
Deep learning approaches: A common pipeline for DL-based audio-visual ADA starts with learning short-term/frame-level behaviours, which are then encoded as clip-level representations to make the clip-level multi-modal depression assessment. For example, Uddin et al. [258] first split the audio-visual data into a set of fixed-length segments and then feed them to a spatio-temporal network for depression-related segment-level multi-modal feature learning, identifying behaviours that are mostly contributed to depression assessment. These segment-level features are finally aggregated as clip-level representation for ADA. Tao et al. [248] separately encode the audio and visual segment-level features as clip-level representations via a temporal attentive pooling, which are then fused via a multi-modal factorised bilinear pooling to provide the final prediction. Niu et al. [29] propose a spatio-temporal attention network to separately learn short-term facial and audio depression cues, and then apply the eigen-evolution pooling to aggregate all short-term audio-visual cues as the clip-level representation. Finally, an attention module and a SVR complementarily combines clip-level multi-modal representations for ADA. Similarly, Yoon et al. [68] employ a two-stream transformer to learn clip-level facial and audio depression representations from all frame-level facial landmarks and audio LLDs, respectively. Then, a cross-attention module is adopted to capture the relationship between audio and visual features to learn a clip-level multi-modal representation for ADA. Focusing on fair ADA, Cheong et al. [244] propose a causal multi-modal framework for debiasing multi-modal behaviours expressed by individuals with different backgrounds where audio and visual features are adaptively fused via a novel referee-based individual fairness guided fusion mechanism to make the multi-modal depression prediction.
@@
Hybrid approaches: Some approaches also manage to leverage both domain knowledge provided by hand-crafted features and task-specific feature learning capability of the DL models for ADA. The AVEC 2019 challenge baseline [63] summarise various audio LLDs and facial behaviour primitives (e.g., AUs) of the given audio-visual clip. Jan et al. [265] extract multiple hand-crafted frame-level facial features (LBP, EOH and LPQ), with a CNN learning depression-specific frame-level facial features. Meanwhile, spectral audio LLDs and MFCC features are extracted from every audio segment. The dynamics of these audio and visual features are then modelled by the feature dynamic history histogram (FDHH) for clip-level ADA. Hu et al. [259] propose a two-stream CNN-LSTM network which separately capture short-term dynamics/features from all frame-level facial primitives (i.e., facial landmarks, gaze and AUs), and audio log-mel spectrograms. Then, a spatio-temporal attention pooling is proposed to obtain long-term audio-visual representations, which are fused at multiple temporal scales via a bridge fusion to output the clip-level multi-modal depression prediction.
4.7.3 Audio-visual-language approaches